A) An economy produces three goods: organ, apple and organ juice. It is assumed that half of the organ are bought and consumed as final good; the other half is used to produce organ juice. 


In 2019, 20 KGs of organ are produced at Rs.50/kg, 40 KGs of apple are produced at 100/kg, and 25 Glasses of organ juice are produced at Rs.200/glass.


In 2020, 40 KGs of organ are produced at Rs.80/kg, 60 KGs of apple are produced at 110/kg, and 30 glasses of organ juice are produced at Rs.195/glass.


  Find the following:


(i) Assuming 2019 is the base year; calculate the real and nominal GDP in 2019 and 2020.


B) Analyze how much Pakistan’s GDP and each of its components is affected by the following transactions? Explain your answers


(i) Atiya spends Rs.1000 to purchase dinner at the finest restaurant in Karachi


(ii)Your mother spend weekend cleaning your house


(iii)Dalda Banspati Limited Pakistan builds Rs.50 million worth of cooking oil, but consumers only buy Rs. 45 million worth of them.



1
Expert's answer
2021-01-25T13:51:15-0500

A) Let's calculate the real GDP in 2019 year:

"Real\ GDP_{2019}=0.5\cdot20\ kg\cdot50\ \dfrac{Rs}{kg}+40\ kg\cdot100\ \dfrac{Rs}{kg}+25\ kg\cdot200\ \dfrac{Rs}{kg}=9500\ Rs."

Since 2019 is the base year, the real GDP euals the nominal GDP.


"Real\ GDP_{2019}=Nominal\ GDP_{2019}=9500\ Rs."

Let's calculate the real GDP in 2020 year:

"Real\ GDP_{2020}=0.5\cdot40\ kg\cdot50\ \dfrac{Rs}{kg}+60\ kg\cdot100\ \dfrac{Rs}{kg}+30\ kg\cdot200\ \dfrac{Rs}{kg}=13000\ Rs."

Let's calculate the nominal GDP in 2020 year:

"Nominal\ GDP_{2020}=0.5\cdot40\ kg\cdot80\ \dfrac{Rs}{kg}+60\ kg\cdot110\ \dfrac{Rs}{kg}+30\ kg\cdot195\ \dfrac{Rs}{kg}=14050\ Rs."

B) The GDP can be calculated as follows:


"GDP=C+I+G+(X-M),"

here, "C" is consumer spending, "I" is investment spending, "G" is government spending, "X" is exports of goods and services and "M" is imports of goods and services.

(i) Atiya spends Rs.1000 to purchase dinner at the finest restaurant in Karachi -> it is an example of the consumer spending. So, the consumer spendings increases by 1000 Rs, and all GDP increases by 1000 Rs.

(ii) Your mother spend weekend cleaning your house -> there is nothing buying or consuming, so there is no impact on the GDP.

(iii) Dalda Banspati Limited Pakistan builds Rs.50 million worth of cooking oil, but consumers only buy Rs. 45 million worth of them -> it is an example of consumer spending. So, the consumer spendings increases by 45 million Rs, and all GDP increases by 45 million Rs.